The firm offers a Net 30 customer cash discount of 2/10. Suppose the customer order is...

  1. The firm offers a Net 30 customer cash discount of 2/10. Suppose the customer order is for $1000, and the firm can invest the funds received from the customer on a short-term investment, what is the minimum effective annual rate the investment has to offer for the cash discount to be worthwhile to the firm?
Answer #1

To explain the question first, net 30 means full amount($1000) from customer is due in 30 days.

Cash discount of 2/10 means that customer gets a 2% discount on invoice if the customer pays in full within 10 days.

Assume 'T' is the date on which invoice is raised and T+30 is when the amount is due. Customer gets a 2% discount ($20) if the customer pays in T+10 days

So, if the customer pays $1000 - 2% discount = $980, in T+10 days, and the firm invests the same in some short term investments, value of the short term investment of $980, after T+30 days should be $1000.

So, in short, minimum effective rate of investment is the one that make makes $980 to $1000 in 20 days [ (T+30) - (T-10) ]

The interest earned on $980 in 20 days is $20 or 20/980=2.0408% absolute.

Effective annually will be ((1+2.0408%)^18-1)= 43.86%